OpenStudy (anonymous):

Solve the following system of equations. x - 2y = 6 2x - 4y = 10

Parth (parthkohli):

At the first look, I knew what the solutions are.

Parth (parthkohli):

Hint: Multiply first equation by 2.

OpenStudy (anonymous):

what do u mean

OpenStudy (anonymous):

Isolate either x or y from one of the equations and use this new definition on your variable in the other equation.

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

Looks parallel, eh @Parth?

OpenStudy (anonymous):

If you multiply equation one by 2 you get 2x-4y=12 2x-4y=10 Anything look funny about that?

OpenStudy (anonymous):

x - 2y = 6 2x - 4y = 10 x=6+2y 2(6+2y)=10 x=6+2y 12+4y=10 x=6+2y 4y=-2 x=6+2y y=-1/2 x=7, x=-1/2

OpenStudy (anonymous):

Double check your work @Kathi26 ...

jhonyy9 (jhonyy9):

@Kathi26 how do you think it ?

jhonyy9 (jhonyy9):

@CliffSedge can you solve it please ?

OpenStudy (anonymous):

Sure, like I said above: "If you multiply equation one by 2 you get 2x-4y=12 2x-4y=10" Since 2x-4y=2x-4y, that implies that 12=10, which is false, so there is no solution. The lines are parallel.

jhonyy9 (jhonyy9):

nice thank you

OpenStudy (anonymous):

Method two: x-2y=6 --> y=x/2-3 2x-4y=10 --> y=x/2-5/2 Both lines have the same slope but a different intercept, ergo, parallel, and no solution.

jhonyy9 (jhonyy9):

ok good luck bye

OpenStudy (anonymous):

@nincompoop since the solution to a system of linear equations is where the lines intersect, might as well look at the graph of the lines.